3. A sample of n = 9 scores has a mean of x̄ =13. After one score is added to the sample, the mean is found to be x̄ =12. What is the value of the score that was added?
Answer #1

n=9; mean= 13

Therefore total sum of all scores= 13*9=117

After an addition of a score, new mean = 12

hence total sum of score now = 12*10= 120

therefore the new score added = 120-117

which is equal to 3